THE PERMANENT ELECTRIC DIPOLE MOMENT OF LANTHANUM IMIDE, LaNH

\underline{TIMOTHY C. STEIMLE} and ROBERT R. BOUSQUET, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ, 85287-1604; ANTHONY J. MERER and SCOTT J. RIXON, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, British Columbia V6T 1Z1, Canada.

There are relatively few experimental determinations of permanent electric dipole moments of transition metal containing polyatomic molecules. Here, we report a study of numerous branch features in the B^{2}Σ^{+}, 001 ← X^{2}Σ^{+}, 000 band of LaNH using high-resolution optical Stark spectroscopy. A supersonic molecular beam sample was prepared by using a pulsed laser vaporization/reaction scheme in which lanthanum vapor was reacted with ammonia. Preliminary analysis of the Stark shifts results in μ = 2.2 D for the X^{2}Σ^{+} state and a much smaller value for the B^{2}Σ^{+} state. A comparision with ScNH and the diatomic Group 3 metal oxides will be given.